Project

General

Profile

Actions

Commissioning and Ops Scheduling meeting : 12/1/2016

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Plans for realignment of SSD
      • Survey of TOF0
      • Realignment of EMR
    • DC to contact Health Physics about calibration of neutron monitor whilst we have access
    • Simulation of effect of air in the drift space between tracker and absorber (John N)
  • Upcoming activities : chaired by DC
  • Issues for discussion
    • Faulty water control replacement (John G)
    • Test of repair of upstream tracker leak - requires removal of US PRY (John G)
  • AOB
  • Date of next meeting : January 19th

Minutes (Preliminary at the moment - will be tidied up tomorrow)

Present : Ryan, Ed, Mark Tucker, Paolo, Pierrick, Paul H, John N, John G, Henry N

  • Review of tasks and activities identified at last meeting
    • Plans for realignment of SSD
      • SSD will be realigned before the February run and plans are being put in place to do this. What is needed now
        is a small group of alignment specialists to get together and discuss the procedure. TOF0 has been re-surveyed
        and the EMR has been shifted back to its nominal position.
    • DC to contact Health Physics about calibration of neutron monitor whilst we have access
      • John says that we just need someone to come along with a source to validate the current performance of the monitor,
        rather than a (multi-week) recalibration which would involve sending the monitor to an external service provider.
    • Simulation of effect of air in the drift space between tracker and absorber (John N)
      • JOhn is working on it but has been stymied by a failure of ScotGrid over Christmas. He will present something next week.
  • Plans for the next week
    • Operations calendar is current.
    • AFC will move to R9 on the 18th
    • Henry : got a UPS coming which needs testing. Goes on IOC servers but can't give an accurate schedule yet.
    • Joe is coming down to go help with the fibres. His schedule needs to be better known.
    • Phil Warburton coming down to work with Mark between 19th and 21st of January
  • Other issues
    • Cooling water issue: With MICE fated to run to the end of the year, John was adamant that the current situation
      with the cooling water controls wasn't fit for purpose. A previous request to the MIPO to redo the controls was rejected
      because of worries that the time needed would poison the compressors. He and Colin have a new plan which they wish to
      discuss with the MIPO.
    • Test of upstream tracker leak repair : John has made a series of clamps to go round the patch panel. Installation and
      testing of these clampls will require us to take off the front of the PRY. We should try to fix this problem. This
      activity will be considered for February.
  • AOB
    • Mark Tucker thinks he has localised the leak on Lh2 absorber. Seems to be on the downstream window of the absorber. Will need
      to work on this in R9.

Updated by Boyd, Steven over 7 years ago ยท 23 revisions